# DART: An Automated End-to-End Object Detection Pipeline with Data Diversification, Open-Vocabulary Bounding Box Annotation, Pseudo-Label Review, and Model Training

### v3

@@ 3.3 Open-vocabulary bounding box annotation via GroundingDINO
Grounding DINO is an open-vocabulary object detector that marries a transformer-based detector DINO [5] with grounded pre-training, which can draw bounding boxes for arbitrary objects given inputs as plain text. The key to Grounding DINO’s open-set capability is the tight fusion of language and vision modalities, achieved through three main components: a feature enhancer, a language-guided query selection, and a cross-modality decoder. The feature enhancer fuses language and visual features by interleaving text-to-image and image-to-text cross-attention based on features extracted from text and image backbones in the early stage. The language-guided query selection uses language information (text features) to guide the selection of detection queries from image features, and the cross-modality decoder extends standard cross attention to two layers using image and text features as keys and values right before model outputs.
@@
To address the common phenomenon of multiple object categories in a single image and avoid the subpar performance of concatenating all classes as text prompts, we utilize the information (collected during data preprocessing) on frequently co-occurring object combinations to create another type of prompt dubbed co-occurring prompt. Unlike the original and synonym prompts, where one phrase is presented at a time, the co-occurring prompt concatenates the names of objects commonly appearing together with delimiters into a single text. For instance, three classes of the LP dataset, ”mining trucks,” ”mining excavators,” and ”mining bulldozers,” often appear simultaneously but rarely with other classes. Thus, images categorized into these three classes during preprocessing are given a co-occurring prompt that combines the names of the corresponding classes. This construction rule is necessary because Grounding DINO tends to produce identical boxes but different class labels when the separately given text prompt is semantically similar, which is often the case for co-occurring classes. This issue is significantly alleviated when these semantically similar texts are input into Grounding DINO as one text prompt. However, the ultimate solution to this problem is the class-agnostic NMS, which will be discussed in the next paragraph. Additionally, we apply the same rule for constructing co-occurring prompts for classes with synonyms, iterating through all combinations of these synonyms.
@@
After getting the output bounding boxes and their corresponding confidence scores from original, synonym, and co-occurring prompts for all images, we apply filtering and non-maximum suppression (NMS) to keep only the best bounding boxes. As illustrated in Algorithm 2, we first sift out the ones with confidence scores bigger than 0.5 unless there’s only one box for an image, in which case we take the one regardless. Then, we transform labels derived from various text prompts back to their original class names from synonyms and finally adopt class-agnostic NMS to address the issue of confusing similar labels and keep only the most feasible annotations. The remaining bounding boxes with their given class labels and scores are the candidate pseudo-labels to be reviewed.
@@
While LMMs excel in semantic understanding and comprehension, they face challenges in numeric-related tasks. For instance, directly generating bounding boxes can be problematic, as depicted in a failure case of GPT-4o in 3(a). In contrast, GPT-4o can precisely interpret the bounding box and its associated class label drawn directly on the image as visual hints, demonstrated by its correct judgments of whether the given boxes and labels are both accurate and complete, as shown in (b)-(c) in Figure 3.
@@
Besides the problem statement and requirement specification, the reminder section for the text prompt is also vital for an optimal response. We find that GPT-4 models tend to be overly strict for bounding box edges and inconsistent when dealing with occluded objects. Consequently, we summarize these common issues in text form (see LABEL:tab:prompt_gpt and other specific requirements for the prompt detailed in Section B.3) and append them at the end of the original text prompt as a reminder. This operation makes GPT-4o much less likely to make the same mistakes. Furthermore, we also utilize Chain-of-Thought (COT) [61] reasoning to enhance the model’s performance. The final text prompt used for pseudo-label review task is detailed in Section B.3.
@@
Like bounding boxes, the generated images via Grouding DINO in Section 3.2 must be validated to ensure they adhere to the requirements of downstream tasks. This verification is crucial as current SDXL models tend to produce overly imaginative images. Compared to the bounding box review based on three criteria, the image photorealism review is an easier task as it only considers photorealism. InternVL-1.5, with its robust visual comprehension, handles this task efficiently. Specifically, it is employed to conduct a preliminary review (with text prompt listed in Table 12) of images generated by SDXL after training with DreamBooth for all selected instances, assessing their photorealism and overall quality. Only the approved images further go through the annotation phase (Section 3.3).
@@ 4.1 Dataset
Data collection. We collect a dataset named Liebherr Products (LP) from the internal database of Libherr, a German-Swiss multinational equipment manufacturer. During data collection, we also employ class and instance identification (in Section 3.1) for bounding box annotation (in Section 3.3) and data diversification (in Section 3.2), respectively.
@@
Dataset composition. The LP dataset (after preprocessing) encompasses more than 15K images of Liebherr’s extensive range of heavy machinery such as construction machines, earthmoving equipment, deep foundation machines, mining machines, various types of cranes, material handling machines, and so on. A list of the exact 23 categories can be found either as the x-axis tick labels in Figure 5 or in the first column of LABEL:tab:synonyms. Figure 5 illustrates the distribution of images in each category, represented by the blue bars labeled ”0:1”. The unbalanced data distribution in the LP dataset makes the downstream object detection task even more challenging. However, as we will see in Figure 9, the final YOLO models still achieve consistently good results across various classes, highlighting the effectiveness of DART. Furthermore, Figure 5 also shows the data distribution after incorporating varying proportions of generated data, which will be elaborated on in the Section 4.5.1.
Dataset split. The dataset is initially split into a training and a test set with an 80/20 distribution. To address class imbalance, stratified random sampling ensures each class is adequately represented in the test set. We allocate 20% of the training set to create a validation set, which is used for hyperparameter fine-tuning. After obtaining the optimal hyperparameter set, the model is trained on the original training with 12K images and evaluated on the 3K images from the test set. The training set could incorporate generated data in specific experiment setups to enhance data diversification. Unless otherwise specified, the generated-to-original data ratio is maintained at the optimal proportion of 3:1, as outlined in Section 4.5.1. Notably, categories that already exhibit strong performance with only the original data are excluded from the augmentation with generated data, as detailed in Section 4.5.4. All performance metrics reported in the tables and figures of this paper are based on the test set.
@@ 4.2 Evaluation metrics
In this work, we utilize the average precision (AP) of the trained object detection models in the last stage as the evaluation metrics for the proposed DART pipeline. AP summarizes the precision-recall (PR) curve into a single value, serving as an approximation of the area under the PR curve. To compute AP, precision $p$ and recall $r$ are first evaluated as follows:
@@ 4.3.1 Implementation of DreamBooth with SDXL
We instantiate DreamBooth with SDXL as the subject-driven image generation framework for the data diversification phase. The SDXL model is preferable, as it significantly outperforms previous versions of Stable Diffusion models such as SD-1.5 [1]. We incorporate prior preservation loss to prevent overfitting and adjust its weight ($\lambda$ in Equation 2) to 1.0. Contrary to SD-1.5, training the text encoder of SDXL has limited effects; we still opt to fine-tune it to enhance overall performance marginally. We extract 69 instances for the training of DreamBooth with SDXL in total. The minimum number of images per instance is set to 3 (as suggested by [9]), while the majority of instances have 10 images each (see Figure 14). We implement dynamic training steps by adjusting the maximum training steps based on the number of images collected for the corresponding instance. 100-160 steps per image are proven to be suitable, combining with the AdamW optimizer [67] with a learning rate $1\times 10^{-4}$ for U-Net and $5\times 10^{-6}$ for the text encoder. The SNR-gamma is set to 5 by default. Other important hyperparameters are explained in Section A.1.
@@ 4.3.2 Implementation of Grounding DINO
Off-the-shelf Grounding DINO base model is employed for bounding box annotation. Through extensive comparative experiments, we determined that a box threshold of 0.27 and a text threshold of 0.25 generally returns high-quality bounding boxes, resulting in only 18 images without annotations, which are consequently excluded from the pipeline. The original collected images and approved generated images are isometrically resized to 800/1333 pixels for the shorter/longer side before going through the model. As for text prompts, we follow instructions given in Section 3.3 to construct all three prompt types as separate inputs to the Grounding DINO model and apply filtering and NMS to the combined results to obtain the candidate bounding boxes to be reviewed. The annotation process is mainly carried out (including speed test as shown in Table 3) by an NVIDIA A100 80GB GPU with FP16 for our case.
@@ 4.4 Ablation study on the overall DART pipeline
Our baseline configuration (#0) utilizes a YOLOv8n model without any additional modules from DART, yielding the lowest performance with an AP of just 0.064. In this setup, the model is trained on one of the most abundant datasets, LVIS [28]. Despite LVIS’s extensive coverage of over 1.2K categories across 164K images, it falls short of capturing all the classes present in our self-collected LP dataset. Consequently, we assign each LP’s category to its closest kin in LVIS during inference, guided by similarity scores calculated from class names and descriptions by SBERT [68]. The results, however, are still disastrous, with a merely non-zero AP of 0.064, suggesting the necessity of the DART pipeline. Integrating the bounding box annotation stage of DART improves the AP to a usable level 0.771. Further incorporation of GPT-4o to rule out inferior pseudo-labels shows an incremental performance increase, resulting in a proper AP of 0.802 and thereby indicating the additional value of LMMs for the DART pipeline. The full power of DART is achieved by the final integration of data diversification. Adding data generated by SDXLs trained for each instance in a DreamBooth framework (and approved by InternVL-1.5) to the original training set boosts the final best AP to 0.832 for the YOLOv8n model, underlining the superior efficacy of the comprehensive DART pipeline.
@@ 4.5.1 Analysis of DreamBooth with SDXL
Table 2 presents the results of training YOLOv8 models with different ratios of generated to original data. When the training data comprises entirely generated data (0:1 ratio), the YOLOv8n model achieves an AP of only 0.031. This outcome is expected due to two primary factors. First, our objective of data diversification drives us to intentionally describe scenarios in the text prompts during inference that are entirely different from the original images (see Section B.2). Secondly, the current DreamBooth with SDXL cannot produce results identical to the original objects but with slight differences. Both factors contribute to a significant disparity between the train and test data distributions. In other words, we deliberately induce an extreme data shift, leading to poor performance on the test set. Conversely, using only original data (ratio 0:1) and hereby eliminating both issues boosts the AP to 0.815.
@@
However, as we gradually increase the amount of generated data, the AP continues to rise, reaching 0.832 for YOLOv8n at a 3:1 ratio (see also Figure 6 for a visualization of this increase). This phenomenon is due to the influence of another opposing mechanism to data shift: introducing diversified data simultaneously enriches the training data diversity, benefiting the model’s learning process. The changes in model performance at different ratios result from the interplay between data shift and enrichment mechanisms. As shown in the Figure 6, further increasing the generated data causes the data shift mechanism to slightly dominate, adversely affecting test performance.
@@
In summary, original data is indispensable for a decent test performance, while generated data offers significant benefits, especially when at a ratio when data shift and diversification mechanisms balance. For the LP dataset, the optimal scenario involves using three times the generated data as the original data. The almost identical trends observed in both YOLOv8n and YOLOv8s models clearly illustrated in Figure 6 corroborate this conclusion.
@@ 4.5.2 Analysis of Grounding DINO
Two Grounding DINO variants, which utilize Swin tiny and base [69] as the backbone, respectively, are tested for the bounding box annotation task. It can be seen from Table 3 that using the more powerful base variant increases test performance for the YOLOv8n model, suggesting better consistency in bounding box annotations and label assignment. However, both of the Grounding DINO variants run too slow to fit the real-time requirements, even on one of the most powerful GPUs, NVIDIA A100. Consequently, we only adopt the Grounding DINO base model for all data annotation tasks for better bounding box quality despite a slightly slower speed, which is negligible for the offline labeling process.
Table 3: Comparsions between Grounidng DINO tiny and base.
@@ 4.5.3 Analysis of LMM
Figure 8: Comparisions of bounding box review consistency. The left confusion matrix displays the agreement between GPT-4o and the human average, whereas the right matrix compares the annotations of two experienced human labelers. GPT-4o demonstrates human-level performance in bounding box review given the same text and visual information.
@@
We design a scenario to validate the effectiveness of both LMMs by comparing the test performance of a YOLOv8n model trained with and without the images they disapproved. First, we construct a dataset using the optimal generated-to-original data ratio of 3:1 obtained in a Section 4.5.2. As indicated in the first row of Table 5, training a YOLOv8n directly on this dataset results in a test performance of AP 0.782. Subsequently, when we remove the images not approved by GPT-4o from the original dataset and retrain the model, the AP climbs up to 0.818. Note that this is already better than training without any generated data, which provides an AP of 0.802 as reported in case #2 of Table 1, exhibiting not only the clear advantages of employing GPT-4o-based filtering but also the pure benefit of data diversification. Finally, by applying InternVL-1.5 to sift through images generated by SDXL and retain the photorealistic ones, the trained model achieves the best performance with an AP of 0.832.

### v4

@@ 3.3 Open-vocabulary bounding box annotation via GroundingDINO
Grounding DINO is an open-vocabulary object detector that marries a transformer-based detector DINO [5] with grounded pre-training, which can draw bounding boxes for arbitrary objects given inputs as plain text. The key to Grounding DINO’s open-set capability is the tight fusion of language and vision modalities, achieved through three main components: a feature enhancer, a language-guided query selection, and a cross-modality decoder. The feature enhancer fuses language and visual features by interleaving text-to-image and image-to-text cross-attention based on features extracted from text and image backbones in the early stage. The language-guided query selection uses language information (text features) to guide the selection of detection queries from image features, and the cross-modality decoder extends standard cross-attention to two layers using image and text features as keys and values right before model outputs.
@@
To address the common phenomenon of multiple object categories in a single image and avoid the subpar performance of concatenating all classes as text prompts, we utilize the information (collected during data preprocessing) on frequently co-occurring object combinations to create another type of prompt dubbed ”co-occurring prompt”. Unlike the original and synonym prompts, where one phrase is presented at a time, the co-occurring prompt concatenates the names of objects commonly appearing together with delimiters into a single text. For instance, three classes of the LP dataset, ”mining trucks,” ”mining excavators,” and ”mining bulldozers,” often appear simultaneously but rarely with other classes. Thus, images categorized into these three classes during preprocessing are given a co-occurring prompt that combines the names of the corresponding classes. This construction rule is necessary because Grounding DINO tends to produce identical boxes but different class labels when the separately given text prompt is semantically similar, which is often the case for co-occurring classes. This issue is significantly alleviated when these semantically similar texts are input into Grounding DINO as one text prompt. However, the ultimate solution to this problem is the class-agnostic NMS, which will be discussed in the next paragraph. Additionally, we apply the same rule for constructing co-occurring prompts for classes with synonyms, iterating through all combinations of these synonyms.
@@
After getting the output bounding boxes and their corresponding confidence scores from the original, synonym, and co-occurring prompts for all images, we apply filtering and non-maximum suppression (NMS) to keep only the best bounding boxes. As illustrated in Algorithm 2, we first sift out the ones with confidence scores bigger than 0.5 unless there’s only one box for an image, in which case we take the one regardless. Then, we transform labels derived from various text prompts back to their original class names from synonyms and finally adopt class-agnostic NMS to address the issue of confusing similar labels and keep only the most feasible annotations. The remaining bounding boxes with their given class labels and scores are the candidate pseudo-labels to be reviewed.
@@
While LMMs excel in semantic understanding and comprehension, they still struggle with tasks that require precise numerical reasoning or manipulation. For instance, directly generating bounding boxes can be problematic, as depicted in a failure case of GPT-4o in 3(a). In contrast, GPT-4o can precisely interpret the bounding box and its associated class label drawn directly on the image as visual hints, demonstrated by its correct judgments of whether the given boxes and labels are both accurate and complete, as shown in (b)-(c) in Figure 3.
@@
Besides the problem statement and requirement specification, the reminder section for the text prompt is also vital for an optimal response. We find that GPT-4 models tend to be overly strict for bounding box edges and inconsistent when dealing with occluded objects. Consequently, we summarize these common issues in text form (see LABEL:tab:prompt_gpt and other specific requirements for the prompt detailed in Section B.3) and append them at the end of the original text prompt as a reminder. This operation makes GPT-4o much less likely to make the same mistakes. Furthermore, we also utilize Chain-of-Thought (COT) [61] reasoning to enhance the model’s performance. The final text prompt used for the pseudo-label review task is detailed in Section B.3.
@@
Like bounding boxes, the generated images via Grounding DINO in Section 3.2 must be validated to ensure they adhere to the requirements of downstream tasks. This verification is crucial as current SDXL models tend to produce overly imaginative images. Compared to the bounding box review based on three criteria, the image photorealism review is an easier task as it only considers photorealism. InternVL-1.5, with its robust visual comprehension, handles this task efficiently. Specifically, it is employed to conduct a preliminary review (with text prompt listed in Table 12) of images generated by SDXL after training with DreamBooth for all selected instances, assessing their photorealism and overall quality. Only the approved images further go through the annotation phase (Section 3.3).
@@ 4.1 Dataset
Data collection. We collect a dataset named Liebherr Products (LP) from the internal database of Liebherr, a German-Swiss multinational equipment manufacturer. During data collection, we also employ class and instance identification (in Section 3.1) for bounding box annotation (in Section 3.3) and data diversification (in Section 3.2), respectively.
@@
Dataset composition. The LP dataset (after preprocessing) encompasses more than 15K images of Liebherr’s extensive range of heavy machinery, such as construction machines, earthmoving equipment, deep foundation machines, mining machines, various types of cranes, material handling machines, and so on. A list of the exact 23 categories can be found either as the x-axis tick labels in Figure 5 or in the first column of LABEL:tab:synonyms. Figure 5 illustrates the distribution of images in each category, represented by the blue bars labeled ”0:1”. The unbalanced data distribution in the LP dataset makes the downstream object detection task even more challenging. However, as we will see in Figure 9, the final YOLO models still achieve consistently good results across various classes, highlighting the effectiveness of DART. Furthermore, Figure 5 also shows the data distribution after incorporating varying proportions of generated data, which will be elaborated on in the Section 4.5.1.
Dataset split. The dataset is initially split into a training and a test set with an 80/20 distribution. To address class imbalance, stratified random sampling ensures each class is adequately represented in the test set. We allocate 20% of the training set to create a validation set, which is used for hyperparameter fine-tuning. After obtaining the optimal hyperparameter set, the model is trained on the original training set with 12K images and evaluated on the 3K images from the test set. The training set could incorporate generated data in specific experiment setups to enhance data diversification. Unless otherwise specified, the generated-to-original data ratio is maintained at the optimal proportion of 3:1, as outlined in Section 4.5.1. Notably, categories that already exhibit strong performance with only the original data are excluded from the augmentation with generated data, as detailed in Section 4.5.4. All performance metrics reported in the tables and figures of this paper are based on the test set.
@@ 4.2 Evaluation metrics
In this work, we utilize the average precision (AP) of the trained object detection models in the last stage as the evaluation metric for the proposed DART pipeline. AP summarizes the precision-recall (PR) curve into a single value, serving as an approximation of the area under the PR curve. To compute AP, precision $p$ and recall $r$ are first evaluated as follows:
@@ 4.3.1 Implementation of DreamBooth with SDXL
We instantiate DreamBooth with SDXL as the subject-driven image generation framework for the data diversification phase. The SDXL model is preferable, as it significantly outperforms previous versions of Stable Diffusion models such as SD-1.5 [1]. We incorporate prior preservation loss to prevent overfitting and adjust its weight ($\lambda$ in Equation 2) to 1.0. Contrary to SD-1.5, training the text encoder of SDXL has limited effects; we still opt to fine-tune it to enhance the overall performance marginally. We extract 69 instances for the training of DreamBooth with SDXL in total. The minimum number of images per instance is set to 3 (as suggested by [9]), while the majority of instances have 10 images each (see Figure 14). We implement dynamic training steps by adjusting the maximum training steps based on the number of images collected for the corresponding instance. 100-160 steps per image are proven to be suitable, combining with the AdamW optimizer [67] with a learning rate $1\times 10^{-4}$ for U-Net and $5\times 10^{-6}$ for the text encoder. The SNR-gamma is set to 5 by default. Other important hyperparameters are explained in Section A.1.
@@ 4.3.2 Implementation of Grounding DINO
The off-the-shelf Grounding DINO base model is employed for bounding box annotation. Through extensive comparative experiments, we determined that a box threshold of 0.27 and a text threshold of 0.25 generally return high-quality bounding boxes, resulting in only 18 images without annotations, which are consequently excluded from the pipeline. The original collected images and approved generated images are isometrically resized to 800/1333 pixels for the shorter/longer side before going through the model. As for text prompts, we follow instructions given in Section 3.3 to construct all three prompt types as separate inputs to the Grounding DINO model and apply filtering and NMS to the combined results to obtain the candidate bounding boxes to be reviewed. The annotation process is mainly carried out (including speed test as shown in Table 3) by an NVIDIA A100 80GB GPU with FP16 for our case.
@@ 4.4 Ablation study on the overall DART pipeline
Our baseline configuration (#0) utilizes a YOLOv8n model without any additional modules from DART, yielding the lowest performance with an AP of just 0.064. In this setup, the model is trained on one of the most abundant datasets, LVIS [28]. Despite LVIS’s extensive coverage of over 1.2K categories across 164K images, it falls short of capturing all the classes present in our self-collected LP dataset. Consequently, we assign each LP’s category to its closest kin in LVIS during inference, guided by similarity scores calculated from class names and descriptions by SBERT [68]. The results, however, are still disastrous, with a merely non-zero AP of 0.064, suggesting the necessity of the DART pipeline. Integrating the bounding box annotation stage of DART improves the AP to a usable level of 0.771. Further incorporation of GPT-4o to rule out inferior pseudo-labels shows an incremental performance increase, resulting in a proper AP of 0.802 and thereby indicating the additional value of LMMs for the DART pipeline. The full power of DART is achieved by the final integration of data diversification. Adding data generated by SDXLs trained for each instance in a DreamBooth framework (and approved by InternVL-1.5) to the original training set boosts the final best AP to 0.832 for the YOLOv8n model, underlining the superior efficacy of the comprehensive DART pipeline.
@@ 4.5.1 Analysis of DreamBooth with SDXL
Table 2 presents the results of training YOLOv8 models under varying ratios of generated to original training data, highlighting the influence of data shift and diversification. When trained exclusively on generated data (1:0 ratio), the YOLOv8n model achieves a low AP of 0.031. This poor performance stems from a significant distributional mismatch between the original and generated datasets by design. Specifically, our prompt design for generation emphasizes maximal novelty to ensure semantic diversity (see Section B.2), which deliberately deviates from the distribution of the original images. Moreover, although DreamBooth with SDXL can produce images with fine-grained resemblance to the original objects, it does not consistently preserve identity-level details. Together, these factors induce an extreme data shift, severely impairing generalization to the original test distribution. By contrast, training solely on original data (0:1 ratio) eliminates this shift and yields an AP of 0.815.
@@
Interestingly, as the proportion of generated data increases alongside original data, model performance improves rather than degrades, revealing the effect of data diversification. At a 3:1 ratio (75 % of the data are generated), the YOLOv8n model achieves its highest AP of 0.832 (see Figure 6). This performance gain arises from the enriched diversity in training samples, which promotes broader feature learning and improved generalization. The overall trend thus reflects a trade-off: while excessive data shift alone is detrimental, a moderate inclusion of diverse synthetic data can counterbalance this by expanding the model’s exposure to varied contexts. However, beyond a certain threshold, the harmful effects of data shift begin to outweigh the benefits of a marginal increase in diversification, leading to a slight decline in AP.
@@
In summary, original data is indispensable for a decent test performance, while generated data offers significant benefits, especially when the ratio is such that the impact of data shift and diversification mechanisms balance. For the LP dataset, the optimal scenario involves using three times the generated data as the original data. The almost identical trends observed for both YOLOv8n and YOLOv8s in Figure 6 further corroborate this conclusion.
@@ 4.5.2 Analysis of Grounding DINO
Two Grounding DINO variants, which utilize Swin tiny and base [69] as the backbone, respectively, are tested for the bounding box annotation task. It can be seen from Table 3 that using the more powerful base variant increases test performance for the YOLOv8n model, suggesting better consistency in bounding box annotations and label assignment. However, both of the Grounding DINO variants run too slow to fit the real-time requirements, even on one of the most powerful GPUs, NVIDIA A100. Consequently, we only adopt the Grounding DINO base model for all data annotation tasks for better bounding box quality, despite a slightly slower speed, which is negligible for the offline labeling process.
Table 3: Comparisons between Grounidng DINO tiny and base.
@@ 4.5.3 Analysis of LMM
Figure 8: Comparisons of bounding box review consistency. The left confusion matrix displays the agreement between GPT-4o and the human average, whereas the right matrix compares the annotations of two experienced human labelers. GPT-4o demonstrates human-level performance in bounding box review given the same text and visual information.
@@
We design a scenario to evaluate the effectiveness of both LMMs by comparing the test performance of a YOLOv8n model trained with and without the images they disapproved. First, we construct a dataset using the optimal generated-to-original data ratio of 3:1, as identified in Section 4.5.2. As shown in the first row of Table 5, training a YOLOv8n model directly on this unfiltered dataset results in an AP of 0.782. In contrast, training solely on the original data achieves a higher AP of 0.802, indicating that naively adding generated images—without any quality filtering—can degrade performance. This underscores the necessity of effective filtering: when the same dataset is curated using GPT-4o to remove low-quality generated samples, the AP rises to 0.818. This not only demonstrates the advantage of GPT-4o-based filtering but also confirms the benefit of data diversification when guided by reliable quality control. Finally, applying InternVL-1.5 to filter for photorealism among SDXL-generated images leads to the best performance, achieving an AP of 0.832.